A ladder 20 meters long leans against a wall. If the bottom of the ladder is being pushed towards the wall at the rate of 20 m/min, how fast is the top of the ladder moving up the wall when the top of the ladder is 6 meters from the ground?

1
Expert's answer
2022-06-06T16:45:29-0400

Let "AB"  be the ladder, where "AB=20 m." Let at time "t"  minutes, the end "A"  of the ladder be "x"  metres from the wall and the end "B"  be "y"  metres from the ground.



Since, "AOB"  is a right angled triangle, by Pythagoras theorem.


"x^2+y^2=20^2"

Differentiate both sides wth respect to "t"


"2x(\dfrac{dx}{dt})+2y(\dfrac{dy}{dt})=0"

Solve for "\dfrac{dy}{dt}"


"\dfrac{dy}{dt}=-\dfrac{x}{y}(\dfrac{dy}{dt})"

Given "\dfrac{dx}{dt}=-20m/min, y=6m."

Substitute


"x=\sqrt{20^2-6^2}=2\sqrt{91}"

"\dfrac{dy}{dt}=-\dfrac{2\sqrt{91}}{6}(-20m/min)=\dfrac{20\sqrt{91}}{3}m/min"

The top of the ladder is sliding up the wall, at the rate of  "\dfrac{20\sqrt{91}}{3}m/min."